# VisJudge-Bench: Aesthetics and Quality Assessment of Visualizations

### v2

@@ 5.2.5 Can VisJudge Generalize to Real-world Applications?
Visualization Generation. VisJudge provides feedback to MatPlotAgent (Yang et al., 2024) for iterative quality improvement. Table 7 shows consistent improvements across seven generation models (+6.07 points average). Notably, base models without domain fine-tuning often degrade performance (7B: -2.39, 72B: -0.61 average). This degradation stems from their systematic evaluation biases (Section 5.2): score inflation and poor discriminative capability lead to misleading feedback that misdirects generation models toward suboptimal outputs. In contrast, VisJudge’s accurate quality assessment consistently improves all generators. Even state-of-the-art models like GPT-5 and Claude-4-Sonnet achieve +4.92 and +7.73 improvements, demonstrating that domain-specific fine-tuning is essential for providing effective feedback.
Visualization Recommendation. VisJudge as a reward model in HAIChart (Xie et al., 2024) improves recommendation accuracy on VizML dataset: +4.40% (Data Queries Hit@1), +2.50% (Overall Hit@1), +5.30% (Overall Hit@3). While 7B and 72B base models show minimal to moderate improvements, VisJudge achieves consistently larger gains across all metrics, outperforming both base models and the baseline (detailed results in Appendix E.1). This validates VisJudge’s broad applicability across generation and recommendation tasks.
(a) Training data scale vs. human-model correlation
(b) Training data scale vs. prediction error
5.3 How Does Training Data Scale Affect Model Performance?
@@ 7 Limitations and Future Work
While VisJudge-Bench demonstrates significant value in visualization quality assessment, it has inherent limitations. The dataset primarily focuses on static visualizations with limited coverage of dynamic evolution and interactive exploration. Additionally, since samples are mainly collected from the web where raw data is often unavailable, our evaluation framework aligns with real-world human assessment by primarily relying on visual presentation. Future work includes expanding dynamic visualization samples with temporal evaluation dimensions, exploring channels to acquire samples with raw data for more comprehensive data fidelity verification, and investigating distribution-based evaluation methods to better capture the diversity of human preferences. To facilitate reproducible research and community development, we will publicly release VisJudge-Bench along with all annotations, quality control records, and the trained VisJudge model weights. We welcome community contributions to improve benchmark quality and coverage.
@@
B.6  Error Analysis of Multi Visualizations and Dashboard ........................................................................................................................................................................B.6
@@
E.1  Generalization to Real-World Applications ........................................................................................................................................................................E.1
E.2  Cross-Architecture Generalization ........................................................................................................................................................................E.1.3
@@ A.1 Data Collection and Construction Pipeline
  • Heuristic Content Pre-screening: We conducted a rapid pre-assessment of image content using programmatic analysis techniques. By employing an edge detection algorithm (ImageFilter.FIND_EDGES) and color complexity analysis (counting the number of unique colors), we discarded a significant number of images that were either too simple (e.g., solid-color backgrounds, blank images) or too complex (e.g., real-world photographs), as these typically do not represent data visualizations.
@@ A.2.2 Visualization Type and Subtype Analysis
For the single visualizations (1,041 samples, 22 subtypes), we observe a clear head-tail distribution. Four basic chart types, bar, pie, line, and area charts—account for 46.1% (480 samples), reflecting their central role in practical visualization design. A second tier of commonly used analytic charts (e.g., treemap, Sankey, heatmap, scatter, histogram, funnel, bubble) contributes 38.0% (396 samples). The remaining 15.9% (165 samples) consists of more specialized forms, where network graphs make up about 2.2% and map-based views (choropleth and point maps) around 3.6%, with the rest spread across other rare, domain-specific encodings. This skewed distribution is in line with  (Borkin et al., 2013) empirical studies of real-world visualization usage.
@@
Among the dashboard samples (1,025 samples, 5 subtypes), Analytical dashboards are overwhelmingly dominant, constituting 72.5% (743 samples) of the dataset. The remainder consists of Operational (11.9%), Interactive (8.9%), and Strategic (6.0%) types, with a negligible fraction (0.7%) classified as other forms. This distribution, heavily weighted towards analysis and decision oriented tools, mirrors the real-world prevalence observed in genre analyses by  (Sarikaya et al., 2019) and design surveys by  (Bach et al., 2023).
@@ A.2.4 Application Domain Distribution Analysis
We annotated the dataset using high-level categories from the IPTC taxonomy ( (Li et al., 2025b)). About 6.2% of samples were grouped as "other" due to limited context. Since the dataset is constructed from real-world web sources through our systematic crawling pipeline (see Appendix A.1), its domain distribution naturally reflects actual visualization demand and usage patterns. The annotated subset shows a clear concentration: economy, business and finance (26.3%), science and technology (20.3%), and labor (15.7%) together account for 62.3% of all labels, aligning with high practical demand in these fields. A middle tier includes society (6.6%) and domains like lifestyle and leisure, health, and politics and government (ranging from 2.7% to 4.9%), while topics like crime, law and justice and religion are rare (<1%). This concentration is even more pronounced in dashboards, where the top three categories comprise 78.5% of the data, reflecting the dominant role of business analytics in dashboard applications.
@@ A.2.6 Low-Quality Sample Distribution and Design Flaw Analysis
Single visualizations. Quality issues in this subset are unevenly distributed, clustering heavily in complex or less common categories. Five specific categories, Treemaps, Scatter Plots, Network Graphs, Heatmaps, and Pie Charts, collectively accountfor 57.7% of all low-quality instances. Across dimensions, design style emerges as the primary bottleneck (mean 2.623, 34% low-score rate), followed by semantic readability (2.831, 31%) and insight discovery (2.779, 32%). Our analysis reveals that the mechanisms of failure diverge significantly based on chart complexity. For complex chart types, quality problems are typically multi-dimensional rather than isolated to a single criterion. In Candlestick Chart, 55% of such samples fail on at least three dimensions simultaneously, as semantic illegibility triggers a chain reaction that degrades data fidelity and insight discovery. Conversely, standard formats like Bar and Line charts generally maintain robust fidelity and readability (mean approx 3.000) while suffering primarily from isolated lacks in aesthetic refinement rather than fundamental communicative breakdowns.
Multiple visualizations. While Comparison Views have a low-score rate (7.2%), their predominance in the dataset means they contribute over half (51.1%) of all low-quality multi-views. A critical finding here is the divergence between visual and semantic performance. Visual composition and color harmony achieve high stability (mean > 3.150, low-score rates < 12%), yet semantic readability and insight discovery lag significantly (mean < 2.700, low-score rates > 33%). This indicates that the core challenge in multi-view design is not layout or coloring, but semantic integration—users struggle to bridge information across views. Furthermore, these semantic failures are strongly coupled with Data Fidelity: in Comparison Views, approximatlly 20% of cases fail on data fidelity, semantic readability, and insight discovery simultaneously, pointing to a systemic inability to convey trustworthy information.
@@
To illustrate these design flaw patterns, we provide detailed case studies in Appendix B: Appendix  B.1,  B.2, and  B.3 present representative cases across high-score (>4), medium-score (2-4), and low-score (<2) ranges for different visualization types; Appendix  B.4 provides typical low-score examples and problem analyses for the three core dimensions—data fidelity, expressiveness, and aesthetics. These cases clearly demonstrate common quality issues and failure modes for each visualization type.
Figure 12: Quality score distributions across visulization subtype.
@@ A.3.2 Annotation Task Design
Our VisJudge-Bench contains 3,090 visualizations, each requiring evaluation across 6 dimensions. To ensure annotation quality and allow annotators to familiarize themselves with the evaluation criteria, we organized the annotations into batches of 15 images per task, with each batch carefully balanced to include 5 single visualizations, 5 multiple visualizations, and 5 dashboards. Before starting each task, annotators were presented with detailed explanations of the evaluation framework, including the meaning and significance of each dimension (Fidelity, Expressiveness, and Aesthetics), enabling them to quickly understand the task requirements and evaluation standards. With 6 questions per image, each annotation task comprised 90 questions and typically took 30–60 minutes, depending on annotator familiarity with the criteria and visualization complexity. Each task was independently annotated by three qualified participants to ensure reliability through majority voting and enable inter-annotator agreement analysis. Based on an estimated hourly wage of \$10 USD, this process ensured high-quality data collection.
@@ A.3.4 Crowdsourcing Quality Control and Candidate Strategy Generation
Stage 2: Candidate Strategy Generation for Expert Review. In the expert adjudication stage, we further designed a systematic conflict identification and resolution mechanism based on representative studies in crowdsourcing quality control, statistical outlier detection, and multi-dimensional evaluation theory (Gadiraju et al., 2015; Rousseeuw and Leroy, 2005; Brennan, 2001). This mechanism provides algorithmic candidate strategies to serve as reference signals for expert review, without replacing expert judgment.
@@ A.4.1 Overall Annotation Quality Assessment
Notably, the crowd-expert MAE was lower than the corresponding crowd-crowd MAE for all three visualization types. This indicates that the deviation between each annotator and the expert standard was actually smaller than the deviations among annotators themselves, suggesting that expert ratings have good representativeness and can effectively balance different annotator perspectives. For the 1–5 rating scale used in this study, our observed MAE values (crowd-crowd: 0.64–0.70, crowd-expert: 0.54–0.59) are comparable to acceptable levels reported in prior crowdsourced visualization studies (Lan and Liu, 2024a) and subjective evaluation task studies (Li et al., 2025b), indicating that annotation quality falls within a reasonable range.
@@ A.4.4 Dimension-Specific Quality Analysis
Single View (159 samples, 31.1%): Data fidelity was adjusted most frequently (20.06%), followed by semantic readability (18.58%), while aesthetics-related dimensions such as color harmony (14.75%), design style (15.34%), and insight discovery (15.04%) had relatively lower adjustment rates. This indicates that single views are most sensitive to data truthfulness and semantic interpretation.
Multi-View (153 samples, 29.9%): Semantic readability had the highest adjustment rate (20.46%), followed by design style (19.02%) and data fidelity (18.73%), while color harmony had the lowest adjustment rate (11.24%). This suggests that multi-view scenarios show the largest disagreements in semantic communication and overall design, but relatively fewer color coordination issues.
@@ B.5 Model Error Analysis Cases
B.6 Error Analysis of Multi Visualizations and Dashboard
@@
Our evaluation identifies a distinct negative correlation between visualization complexity and model performance. As visualization complexity increases, from single charts to multi-view compositions and dashboards, the efficacy of current multimodal models notably declines. In the analyzed subset comprising 139 multi visualization sets and 94 dashboard cases, quality scores on visual composition and color harmony drop sharply. In dashboards, this degradation extends beyond aesthetic metrics, significantly impacting in insight discovery and semantic readability. Based on our analysis, we attribute that degradation is the lack of cross-view visual understanding. The models tend to process each view as an isolated visualization and rely heavily on local visual cues (e.g., grid alignment or color consistency) while failing to construct the cross-view semantic connections and analytical pathways characteristic of human expert interpretation.
@@
Visual Composition and Color Harmony Evaluation. Regarding aesthetic dimensions, the models exhibit a problematic dual bias. They tend to overestimate layout regularity, assigning high scores as long as views are arranged in a standard grid with consistent styling, while disregarding redundancy or the absence of meaningful cross-view relationships (see Figure  6, left). For instance, in the multi visualization subset, this positive bias is widespread: InternVL3, Claude-3.5-Sonnet, and Qwen2.5-VL rated visual composition higher than human experts in roughly 90% of the cases (89.2%–92.1%), inflating scores by over 1.1 points on average. This tendency persists in Dashboards, where Claude-3.5-Sonnet continued to overestimate Color Harmony in 75.5% of cases, inflating the metric by 1.09 points. Simultaneously, they are overly strict about practical design trade-offs. The models tend to enforce strict encoding rules, penalizing visualizations that make necessary compromises, such as minor inconsistencies in color encoding or redundant legends(see Figure  7, right). Gemini-2.5-Pro, for example, strictly penalized color harmony, underestimating scores in 44 multi visualization cases and 29 dashboard cases, with negative deviations nearing 1 point. This indicates that models perform surface-level pattern matching rather than comprehending the functional intent behind design decisions.
@@
Data Fidelity Evaluation. In terms of data fidelity, we find that the models often confuse structural completeness with content validity. Charts containing only axes and frames, but no actual data marks, often receive high fidelity scores merely because they adhere to visualization design standards (Figure  26). This hallucination is pervasive across both open-source and proprietary models. In the multi visualization subset, Qwen2.5-VL and InternVL3 overestimated data fidelity in 95.0% and 96.4% of cases, respectively. Proprietary models like GPT-4o and Claude-3.5-Sonnet followed a similar pattern, inflating fidelity scores in 87.8% of cases with an average increase exceeding 1.50 points. In contrast, simplified schematic views, such as those omitting detailed axis labels to emphasize trend comparison, are frequently deemed unreliable by GPT-5 and Gemini-2.5-Pro due to their deviation from standard forms. This suggests current models prioritize structural conformity to standard charts over the accurate conveyance of underlying data.
@@
Dashboard Insight Discovery and Semantic Readability Evaluation. In dashboards, we observe that models tend to equate interface completeness with analytical depth. Dashboards featuring diverse chart types within a unified visual style often receive high insight discovery and semantic readability scores, even when lacking logical coherence or narrative structure (e.g., Figure  20, top). Qwen2.5-VL exhibited the most extreme positive bias in semantic readability, inflating scores in 90.4% of the Dashboard cases with an average increase of 1.28. Conversely, complex and information-rich dashboards are frequently unfairly penalized by GPT-5 and Gemini-2.5-Pro, which misinterpret their high density as “visual clutter” or “data distortion”. GPT-5 underestimated semantic readability in 33 cases (average decrease of 0.78 points), while Gemini-2.5-Pro showed an even stronger negative bias, lowering scores in 67 cases by an average of 1.15 points. This underscores a critical limitation: current models lack a robust understanding of narrative logic within visualization interfaces and struggle to distinguish between visually polished artifacts and functionally effective analytical tools.
@@
Figure 23: Qwen2.5-VL-7B–Instruct and GPT-5 error cases.

### v3

@@ 5.2.5 Can VisJudge Generalize to Real-world Applications?
Visualization Generation. VisJudge provides feedback to MatPlotAgent (Yang et al., 2024) for iterative quality improvement. Table 4 shows consistent improvements across seven generation models (+6.07 points average). Notably, base models without domain fine-tuning often degrade performance (7B: -2.39, 72B: -0.61 average). This degradation stems from their systematic evaluation biases (Section 5.2.2): score inflation and poor discriminative capability lead to misleading feedback that misdirects generation models toward suboptimal outputs. In contrast, VisJudge’s accurate quality assessment consistently improves all generators. Even state-of-the-art models like GPT-5 and Claude-4-Sonnet achieve +4.92 and +7.73 improvements, demonstrating that domain-specific fine-tuning is essential for providing effective feedback.
Visualization Recommendation. VisJudge as a reward model in HAIChart (Xie et al., 2024) improves recommendation accuracy on the VizML dataset: +4.40% (Data Queries Hit@1), +2.50% (Overall Hit@1), +5.30% (Overall Hit@3). While 7B and 72B base models show minimal to moderate improvements, VisJudge achieves consistently larger gains across all metrics, outperforming both base models and the baseline (detailed results in Appendix E.2). This validates VisJudge’s broad applicability across generation and recommendation tasks.
@@ 7 Limitations and Future Work
VisJudge-Bench has several limitations: (1) it focuses on static visualizations, with limited coverage of dynamic and interactive content; (2) web-collected samples typically lack raw data, so evaluation relies on visual presentation alone, limiting data fidelity verification; and (3) point-estimate scoring may not fully capture the diversity of human preferences. Future work includes expanding dynamic visualization samples with corresponding evaluation criteria, exploring ways to obtain samples with raw data for more thorough fidelity assessment, and investigating distribution-based evaluation methods to better reflect human preference diversity.
@@
Reproducibility Statement
To ensure reproducibility, we provide comprehensive documentation and resources. The complete VisJudge-Bench construction process is detailed in Appendix A, covering data collection, filtering, and expert annotation protocols. The six-dimensional evaluation framework and VisJudge implementation are described in Appendices C and D, respectively. We release the complete dataset with 3,090 expert-annotated samples across three categories (single_vis, multi_vis, dashboard) at https://github.com/HKUSTDial/VisJudgeBench, where each sample includes visualization images, six-dimensional scores, and evaluation prompts. All experimental configurations, evaluation metrics (MAE, MSE, Pearson correlation), and human consistency analysis methods are fully documented to enable result reproduction.
Acknowledgments
This paper was supported by the NSF of China (62402409); Youth S&T Talent Support Programme of Guangdong Provincial Association for Science and Technology (SKXRC2025461); the Young Talent Support Project of Guangzhou Association for Science and Technology (QT-2025-001); Guangdong Basic and Applied Basic Research Foundation (2023A1515110545); Guangzhou Basic and Applied Basic Research Foundation (2025A04J3935); and Guangzhou-HKUST(GZ) Joint Funding Program (2025A03J3714).
@@ Acknowledgments
B.6  Error Analysis of Multiple Visualizations and Dashboards ........................................................................................................................................................................B.6
@@
E.1  How Does Training Data Scale Affect Model Performance? ........................................................................................................................................................................E.1
E.2  Generalization to Real-World Applications ........................................................................................................................................................................E.2
E.3  Cross-Architecture Generalization ........................................................................................................................................................................E.2.3
@@ A.1 Data Collection and Construction Pipeline
  • Heuristic Content Pre-screening: We conducted a rapid pre-assessment of image content using programmatic analysis techniques. By employing an edge detection algorithm and color complexity analysis (counting the number of unique colors), we discarded a significant number of images that were either too simple (e.g., solid-color backgrounds, blank images) or too complex (e.g., real-world photographs), as these typically do not represent data visualizations.
@@ A.2.2 Visualization Type and Subtype Analysis
For the single visualizations (1,041 samples, 22 subtypes), we observe a clear head-tail distribution. Four basic chart types—bar, pie, line, and area charts—account for 46.1% (480 samples), reflecting their central role in practical visualization design. A second tier of commonly used analytic charts (e.g., treemap, Sankey, heatmap, scatter, histogram, funnel, bubble) contributes 38.0% (396 samples). The remaining 15.9% (165 samples) consists of more specialized forms, where network graphs make up about 2.2% and map-based views (choropleth and point maps) around 3.6%, with the rest spread across other rare, domain-specific encodings. This skewed distribution is in line with (Borkin et al., 2013) empirical studies of real-world visualization usage.
@@
Among the dashboard samples (1,025 samples, 5 subtypes), Analytical dashboards are overwhelmingly dominant, constituting 72.5% (743 samples) of the dataset. The remainder consists of Operational (11.9%), Interactive (8.9%), and Strategic (6.0%) types, with a negligible fraction (0.7%) classified as other forms. This distribution, heavily weighted towards analysis- and decision-oriented tools, mirrors the real-world prevalence observed in genre analyses by (Sarikaya et al., 2019) and design surveys by (Bach et al., 2023).
@@ A.2.4 Application Domain Distribution Analysis
We annotated the dataset using high-level categories from the IPTC taxonomy (Li et al., 2025c). About 6.2% of samples were grouped as "other" due to limited context. Since the dataset is constructed from real-world web sources through our systematic crawling pipeline (see Appendix A.1), its domain distribution naturally reflects actual visualization demand and usage patterns. The annotated subset shows a clear concentration: economy, business and finance (26.3%), science and technology (20.3%), and labor (15.7%) together account for 62.3% of all labels, aligning with high practical demand in these fields. A middle tier includes society (6.6%) and domains like lifestyle and leisure, health, and politics and government (ranging from 2.7% to 4.9%), while topics like crime, law and justice and religion are rare (<1%). This concentration is even more pronounced in dashboards, where the top three categories comprise 78.5% of the data, reflecting the dominant role of business analytics in dashboard applications.
@@ A.2.6 Low-Quality Sample Distribution and Design Flaw Analysis
Single visualizations. Quality issues in this subset are unevenly distributed, clustering heavily in complex or less common categories. Five specific categories, Treemaps, Scatter Plots, Network Graphs, Heatmaps, and Pie Charts, collectively account for 57.7% of all low-quality instances. Across dimensions, design style emerges as the primary bottleneck (mean 2.623, 34% low-score rate), followed by semantic readability (2.831, 31%) and insight discovery (2.779, 32%). Our analysis reveals that the mechanisms of failure diverge significantly based on chart complexity. For complex chart types, quality problems are typically multi-dimensional rather than isolated to a single criterion. In the Candlestick Chart, 55% of such samples fail on at least three dimensions simultaneously, as semantic illegibility triggers a chain reaction that degrades data fidelity and insight discovery. Conversely, standard formats like Bar and Line charts generally maintain robust fidelity and readability (mean approx 3.000) while suffering primarily from isolated lacks in aesthetic refinement rather than fundamental communicative breakdowns.
Multiple visualizations. While Comparison Views have a low-score rate (7.2%), their predominance in the dataset means they contribute over half (51.1%) of all low-quality multi-views. A critical finding here is the divergence between visual and semantic performance. Visual composition and color harmony achieve high stability (mean > 3.150, low-score rates < 12%), yet semantic readability and insight discovery lag significantly (mean < 2.700, low-score rates > 33%). This indicates that the core challenge in multi-view design is not layout or coloring, but semantic integration—users struggle to bridge information across views. Furthermore, these semantic failures are strongly coupled with Data Fidelity: in Comparison Views, approximately 20% of cases fail on data fidelity, semantic readability, and insight discovery simultaneously, pointing to a systemic inability to convey trustworthy information.
@@
To illustrate these design flaw patterns, we provide detailed case studies in Appendix B: Appendix B.1,  B.2, and  B.3 present representative cases across high-score (>4), medium-score (2–4), and low-score (<2) ranges for different visualization types; Appendix B.4 provides typical low-score examples and problem analyses for the three core dimensions—data fidelity, expressiveness, and aesthetics. These cases clearly demonstrate common quality issues and failure modes for each visualization type.
Figure 11: Quality score distributions across visualization subtypes.
@@ A.3.2 Annotation Task Design
Our VisJudge-Bench contains 3,090 visualizations, each requiring evaluation across 6 dimensions. To ensure annotation quality and allow annotators to familiarize themselves with the evaluation criteria, we organized the annotations into batches of 15 images per task, with each batch carefully balanced to include 5 single visualizations, 5 multiple visualizations, and 5 dashboards. Before starting each task, annotators were presented with detailed explanations of the evaluation framework, including the meaning and significance of each dimension (Fidelity, Expressiveness, and Aesthetics), enabling them to quickly understand the task requirements and evaluation standards. With 6 questions per image, each annotation task comprised 90 questions and typically took 30–60 minutes, depending on annotator familiarity with the criteria and visualization complexity. Each task was independently annotated by three qualified participants to ensure reliability through majority voting and enable inter-annotator agreement analysis. Annotators were compensated at an estimated hourly rate of \$10 USD, which is competitive relative to the platform average and helped attract and retain qualified workers able to devote adequate attention to the task. Together with our strict screening criteria and embedded validation checks, this compensation structure supported the collection of high-quality annotations.
@@ A.3.4 Crowdsourcing Quality Control and Candidate Strategy Generation
Stage 2: Candidate Strategy Generation for Expert Review. In the expert adjudication stage, we further designed a systematic conflict identification and resolution mechanism based on representative studies in crowdsourcing quality control, statistical outlier detection, and multi-dimensional evaluation theory (Gadiraju et al., 2015; Rousseeuw and Leroy, 2003; Brennan, 1992). This mechanism provides algorithmic candidate strategies to serve as reference signals for expert review, without replacing expert judgment.
@@ A.4.1 Overall Annotation Quality Assessment
Notably, the crowd-expert MAE was lower than the corresponding crowd-crowd MAE for all three visualization types. This indicates that the deviation between each annotator and the expert standard was actually smaller than the deviations among annotators themselves, suggesting that expert ratings have good representativeness and can effectively balance different annotator perspectives. For the 1–5 rating scale used in this study, our observed MAE values (crowd-crowd: 0.64–0.70, crowd-expert: 0.54–0.59) are comparable to acceptable levels reported in prior crowdsourced visualization studies (Lan and Liu, 2024) and subjective evaluation task studies (Li et al., 2025c), indicating that annotation quality falls within a reasonable range.
@@ A.4.4 Dimension-Specific Quality Analysis
Single Visualizations (159 samples, 31.1%): Data fidelity was adjusted most frequently (20.06%), followed by semantic readability (18.58%), while aesthetics-related dimensions such as color harmony (14.75%), design style (15.34%), and insight discovery (15.04%) had relatively lower adjustment rates. This indicates that single visualizations are most sensitive to data truthfulness and semantic interpretation.
Multiple Visualizations (153 samples, 29.9%): Semantic readability had the highest adjustment rate (20.46%), followed by design style (19.02%) and data fidelity (18.73%), while color harmony had the lowest adjustment rate (11.24%). This suggests that multiple visualizations show the largest disagreements in semantic communication and overall design, but relatively fewer color coordination issues.
@@ B.5 Model Error Analysis Cases
B.6 Error Analysis of Multiple Visualizations and Dashboards
@@
Our evaluation identifies a distinct negative correlation between visualization complexity and model performance. As visualization complexity increases, from single charts to multi-view compositions and dashboards, the efficacy of current multimodal models notably declines. To conduct an in-depth analysis of model error patterns in complex visualizations, we filtered from the test set those samples where at least one model’s predicted score deviated substantially from the human annotation (absolute deviation >1.0>1.0), yielding a subset of 139 multi-visualization sets and 94 dashboard cases. In this subset, quality scores on visual composition and color harmony drop sharply. In dashboards, this degradation extends beyond aesthetic metrics, significantly impacting insight discovery and semantic readability. Based on our analysis, we attribute this degradation to the lack of cross-view visual understanding. The models tend to process each view as an isolated visualization and rely heavily on local visual cues (e.g., grid alignment or color consistency) while failing to construct the cross-view semantic connections and analytical pathways characteristic of human expert interpretation.
@@
Visual Composition and Color Harmony Evaluation. Regarding aesthetic dimensions, the models exhibit a problematic dual bias. They tend to overestimate layout regularity, assigning high scores as long as views are arranged in a standard grid with consistent styling, while disregarding redundancy or the absence of meaningful cross-view relationships (see Figure 6, left). For instance, in the multi-visualization subset, this positive bias is widespread: InternVL3, Claude-3.5-Sonnet, and Qwen2.5-VL rated visual composition higher than human experts in roughly 90% of the cases (89.2%–92.1%), inflating scores by over 1.1 points on average. This tendency persists in dashboards, where Claude-3.5-Sonnet continued to overestimate Color Harmony in 75.5% of cases, inflating the metric by 1.09 points. Simultaneously, they are overly strict about practical design trade-offs. The models tend to enforce strict encoding rules, penalizing visualizations that make necessary compromises, such as minor inconsistencies in color encoding or redundant legends (see Figure 7, right). Gemini-2.5-Pro, for example, strictly penalized color harmony, underestimating scores in 44 multi-visualization cases and 29 dashboard cases, with negative deviations nearing 1 point. This indicates that models perform surface-level pattern matching rather than comprehending the functional intent behind design decisions.
@@
Data Fidelity Evaluation. In terms of data fidelity, we find that the models often confuse structural completeness with content validity. Charts containing only axes and frames, but no actual data marks, often receive high fidelity scores merely because they adhere to visualization design standards (Figure 25). This hallucination is pervasive across both open-source and proprietary models. In the multi-visualization subset, Qwen2.5-VL and InternVL3 overestimated data fidelity in 95.0% and 96.4% of cases, respectively. Proprietary models like GPT-4o and Claude-3.5-Sonnet followed a similar pattern, inflating fidelity scores in 87.8% of cases with an average increase exceeding 1.50 points. In contrast, simplified schematic views, such as those omitting detailed axis labels to emphasize trend comparison, are frequently deemed unreliable by GPT-5 and Gemini-2.5-Pro due to their deviation from standard forms. This suggests current models prioritize structural conformity to standard charts over the accurate conveyance of underlying data.
@@
Dashboard Insight Discovery and Semantic Readability Evaluation. In dashboards, we observe that models tend to equate interface completeness with analytical depth. Dashboards featuring diverse chart types within a unified visual style often receive high insight discovery and semantic readability scores, even when lacking logical coherence or narrative structure (e.g., Figure 19, top). Qwen2.5-VL exhibited the most extreme positive bias in semantic readability, inflating scores in 90.4% of the dashboard cases with an average increase of 1.28. Conversely, complex and information-rich dashboards are frequently unfairly penalized by GPT-5 and Gemini-2.5-Pro, which misinterpret their high density as “visual clutter” or “data distortion”. GPT-5 underestimated semantic readability in 33 cases (average decrease of 0.78 points), while Gemini-2.5-Pro showed an even stronger negative bias, lowering scores in 67 cases by an average of 1.15 points. This underscores a critical limitation: current models lack a robust understanding of narrative logic within visualization interfaces and struggle to distinguish between visually polished artifacts and functionally effective analytical tools.
@@
Figure 22: Qwen2.5-VL-7B-Instruct and GPT-5 error cases.